OpenStudy (anonymous):

Does (m + 8)^2 equal m^2 + 64?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

(m+8)^2=(m+8)(m+8)=m^2+8m+8m+64=m^2+16m+64

OpenStudy (anonymous):

How come you do it that way? I solve it this way: m^2 + 16?

OpenStudy (anonymous):

m^2+16m+64

OpenStudy (anonymous):

you can do (m+8)^2=(m+8)(m+8)=m^2+8m+8m+64=m^2+16m+64 or you can do like this (m+8)^2 =(m+8)(m+8) =m^2+8m+8m+64 =m^2+16m+64 same answer right ...lol

OpenStudy (anonymous):

is it better OMG that way?